1/4

Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников,

Дмитрий Николаевич Садовников,

Sten’ka Razin Из-за острова на стрежень (Стенька Разин)

Volga, Vol- ga (Stenka Rasin)

Stenka Rasin

tradukita de Tr. Jacob Robbins tradukita de N. N. 20

tradukita de N. N. 21

1. From beyond the wooded island

To the river wide and free Proudly sailed the arrow-

breasted Ships of Cossack yeoman-

ry.

1. Из-за острова на стрежень, На простор речной

волны, Выплывают

расписные, Острогрудые челны.

Bag om øen, på den

brede underskønne Vol-

gaflod Stenjka Rasins stolte

både ud på nye togter

stod.

In den Wellen hinter

Inseln Ziehen Kähne male-

risch, - |: Fangen leis an aufzuwachen, Bang ist jedes Ange-

sicht. :|

2. On the first is Stenka Razin

With a princess at his si-

de, |: Drunken, holds a mar-

riage revel With his beautiful young

bride. :| 2. На переднем Стенька Разин, Обнявшись, сидит с княжной, Свадьбу новую справляет, Сам веселый и хмельной.

På den første høvding Stenjka sidder selv i bådens

stavn, han på kamp og strid ej tænker, har prinsessen i sin

favn.

Stenka Rasin vorn als erster, Selig in der Trunken-

heit, |: Hält im Arme die Prinzessin, Die er eben erst be-

freit. :|

3. But behind them rose a whisper,

“He has left his sword to

woo; |: One short night, and

Stenka Razin Has become a woman

too!” :| 3. А она, потупив

очи, Ни жива и ни мертва,

Молча слушает хмельные Атамановы слова.

Men hans mænd er Donkosakker, de er vrede, for de

ved, han har glemt dem for en kvinde, for en nat i kærlighed.

.................................................. .................................................. .................................................. ..................................................

4. Stenka Razin hears the jeering

Of his discontented band, |: And the lovely Persian

princess He has circled with his

hand. :| 4. Позади их слышен

ропот:

- Нас на бабу променял! Только ночь с ней провозилс Сам наутро бабой

стал .

Stenjka Rasin hører

latter, blodet koger i hans

bryst Over Volgas brede strømme højt nu lyder Stenjkas røst.

.................................................. .................................................. .................................................. ..................................................

…  …  …  …

2/4

5. His black brows have come together

As the waves of anger rise, |: And the blood comes

rushing swiftly To his piercing, jet-black

eyes. :| 5. Этот ропот и насмешки Слышит грозный атаман, И могучею рукою Обнял персиянки

стан.

....................................... ....................................... ....................................... .......................................

.................................................. .................................................. .................................................. ..................................................

6. “I will give you all you ask for,

Life and heart, and head

and hand,” |: Echo rolls the pealing

thunder Of his voice across the

land. :| 6. Брови черные сошлися, Надвигается гроза. Буйной кровью налилис Атамановы глаза.

....................................... ....................................... ....................................... .......................................

.................................................. .................................................. .................................................. ..................................................

7. “Volga, Volga, mother Volga,

Deep and wide beneath

the sun, |: You have never seen a

present From the Cossack of the

Don. :|

7.  ”Ничего  не

пожалею, Буйну голову отдам!“ — Раздается голос

властный По окрестным берегам.

....................................... ....................................... ....................................... .......................................

Plötzlich tönt ein dumpf Gemurre: Er verrät uns um ein

Weib, |: All der Seinen Glück vergißt er Um geringen Zeitvertreib. :|

8. And that peace might rule as always

All my free-born men and

brave, |: Volga, Volga, mother

Volga, Volga, make this girl a

grave.” :| 8. “Волга, Волга, мать родная, Волга, русская река, Не видала ты

подарка От донского казака!

....................................... ....................................... ....................................... .......................................

Wolga, Wolga, Mutter Wolga, Ruhig fließest du da-

hin, |: Ahnst nicht, was ein Donkosake Fürchterliches hat im Sinn. :|

9. With a sudden, mighty movement,

Razin lifts the beauty

high, |: And he casts her where

the waters Of the Volga move and

sigh. :| 9. Чтобы не было раздора Между вольными людьми,

Волга, Волга, мать

родная, На, красавицу возьми!”

“For at ingen splid skal være mellem frie mænd,-

tag så Moder Volga, denne kvinde!” og hans arm om hen-

de lå.

.................................................. .................................................. .................................................. ..................................................

…  …  …  …

3/4

10. Now a silence like the grave sinks

To all those who stand to

see, |: And the battle-hardened

Cossacks Sink to weep on bended

knee. :| 10. Мощным взмахом поднимает Он красавицу

княжну И за борт ее бросает В набежавшую

волну.

Og han ta’r med stærke arme da sin brud og kaster

ned /: i de dybe, mørke

bølger, sine nætters salig-

hed. :/ Vorn als erster Stenka Rasin Hebt das Weib in wilder Wut, - |: Wolga, Wolga, nimm ein Opfer! Und er wirft sie in die Flut. :|

11. “Dance, you fool, and men, make merry!

What has got into your

eyes? |: Let us thunder out a

chanty Of a place where beauty

lies.” :|

11.  “Что  ж

вы, братцы, приуныли? Эй, ты, Филька, черт, пляши! Грянем песню

удалую На помин ее души!..”

“Men hvad fanden, hvorfor hænger I med næbbet? Bliv kun ved! Dans og syng de røversange, syng for hendes sjælefred!”

Und er sieht sie untergehen, Hört noch ihren Jammerschrei: |: Stirb als Opfer meiner Treue! Stenka Rasin, er ist

frei. :|

12. From beyond the wooded island

To the river wide and free, |: Proudly sail the arrow-

breasted Ships of Cossack yeoman-

ry. :| 12. Из-за острова на стрежень, На простор речной

волны, Выплывают

расписные Острогрудые челны.

Frem fra øen - ud på

dybet stævner Stenjka Ra-

sin ud /: med den stærke, stolte både, sejler fra sin døde

brud. :/ Und die Kähne ziehen weiter, Und die Kähne ziehen fort, |: Und die Wolga fließet weiter Über diesen Schreckensort.

…  …  …  …

4/4

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .

Verkinto de tiu ĉi Rusa poemo estas Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883).

Arg-478-927 (2010-03-07 18:59:55)

Pri la aŭtoro vidu ankaŭ la informojn sub http: // ru. wikipedia. org , la ruslingva parto de Wikipedia. Informoj pri la kanto, la melodio kaj diversaj versioj de la kanto estas troveblaj en la retejo http: // a-pesni. golosa. info/ popular20/ izzaostrova. htm .

Traduko de la Ru-

sa  poemo  “Из-
за  острова  на

стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Danan de N. N. 20.

Arg-478-1037 (2010-03-08 12:16:20)

Aperinte en 1947 en la kant-libro “DLEA-Kantaro” eldonita de Dana Laborista Esperanto-Asocio (DLEA), n-ro 100 kaj en la kantlibro “Store Esperanto-Sangbog”, SEFO 1939, n-ro 31. Vidu la retejon esperantonia. dk sub http: // www. esperantonia. dk/ volga. htm .

Traduko de la Ru-

sa  poemo  “Из-
за  острова  на

стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Germanan de N. N. 21.

Arg-478-1038 (2010-03-09 15:31:43)

Mi, Manfred Retzlaff, trovis tiun ĉi germanigon de la kanto en http: // www. musicanet. org/ robokopp/ Lieder/ stenkara. html . Latinskriba aliskribo kaj laŭvorta germanigo troveblas en http: // www. kaikracht. de/ balalaika/ songs/ izza_bal. htm . Vidu ankaŭ la vikipedian retejon http: // de. wikipedia. org/ wiki/ Stenka_Rasin .